OpenStudy (anonymous):

A bird can fly 24km/h . Part A How long does it take to fly 3.1km ? Express your answer to two significant figures and include the appropriate units. Δt =

OpenStudy (theeric):

Hi! I have about 10 minutes, but this is a simple problem after you see what to do. Do you know how to start?

OpenStudy (anonymous):

no I dont I know that there is formula but dont know which to use

OpenStudy (theeric):

Well, first you need to identify what quantities you have! What is 24km/h?

OpenStudy (anonymous):

is that the speed?

OpenStudy (theeric):

Yep! What is 3.1km?

OpenStudy (anonymous):

distance?

OpenStudy (theeric):

Right! And what do you want to find?

OpenStudy (anonymous):

time

OpenStudy (anonymous):

or is it change in time bc its delta T

OpenStudy (theeric):

Right. I agree! :) And you could say it's also the change in distance. (Side note: this is still not a vector.) So you have speed, distance, and need time. You must find or create an equation that has only \(v\), \(d\text{ or }\Delta d\), and \(\Delta t\). What is the definition of speed? Change in distance per time! \(v=\dfrac{\Delta d}{\Delta t}\implies \Delta t=\dfrac{\Delta d}{v}.\)

OpenStudy (theeric):

Now do you see what to do? I must get going. Take care! :)

OpenStudy (anonymous):

ok i will try to figure it out thank you!

OpenStudy (theeric):

Hi! I used the definition of speed, and wrote it in math! Then I used algebra to solve for \(\Delta t\). Now we substitute in our values for speed and distance to see that\[\Delta t= \frac{\Delta d}{v}=\frac{3.1\ [km]}{24\ [km\ /\ hr]}=.1291\bar 6\ [hr]\] So, that is the time! Notice that I included the units. Now, we need to round it to two significant figures. That is, our answer should be \(.13\ [hr]\). Now, that doesn't show what significant figures are. If you're unsure, just do a search of ask! :)

OpenStudy (anonymous):

oki thank you, I actually got that but I forgot to close the question..silly of me...thank you!

OpenStudy (theeric):

I'm glad you got it! You're welcome! :)